OpenStudy (anonymous):

multiple choice algebra 2 pleaseee !

OpenStudy (anonymous):

ya....question????

OpenStudy (anonymous):

x-3y=-6 x+6=3y (x+6)/3=y

OpenStudy (anonymous):

its not showing up the way i type it

OpenStudy (anonymous):

solve for y in the first equation,. x – 3y = –6 2x – 7y = 10 (x + 2, x + 6, –x + 2, x + 2)

OpenStudy (anonymous):

2x+y=11 y=11-2x =-2x+11

OpenStudy (anonymous):

this was second question

OpenStudy (anonymous):

thankyou so much :)

OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x – 3y = –3, 2x – 6y = –6 ( –1, 0 , 1, 2 ,)

OpenStudy (anonymous):

mutiply first equation by 2 and second equation by 5:; 10x-6y=-6............(1) 10x-30y=-30........(2) subtract (2) from (1): 24y=24 y=1

OpenStudy (anonymous):

nice explanation thanks :) do u wanna help me with the other 2 too?

OpenStudy (anonymous):

ok...sure

OpenStudy (anonymous):

this one is kinda tough for me, it involves fractions lol In the below system, solve for y in the first equation. x – 3y = –6 2x – 7y = 10 (x + 2, x + 6 , –x + 2 , x + 2 ,)

OpenStudy (anonymous):

x+6=3y so y=(x+6)/3

OpenStudy (anonymous):

the answer choices are wrong. they are- ( 1/3 x+2, 1/3 x+6, -x+2, -1/3 x+2)

OpenStudy (anonymous):

1/3 x+6 then?

OpenStudy (anonymous):

ya...

OpenStudy (anonymous):

thanks! this one i think its no solutions but im not sure. Use the substitution method to solve the following system of equations. 3x – y = 2 6x – 2y = 4 (1, 1) (–1, –5) No Solutions Infinitely Many Solutions

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

from 1st equation:: y=3x-2 put the value of y in second equation: 6x-2(3x-2)=4 6x-6x+4=4 4=4 which is true so infinitely many solutions

OpenStudy (anonymous):

oh okay thanks for all your help!

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

let me check last on last one

OpenStudy (anonymous):

no its right infintely many solutions.......

OpenStudy (anonymous):

any other

OpenStudy (anonymous):

i have some that are a little more difficult. the have 3 variables. The following system is solved by substitution. Which expression is substituted for z in the second and third equations? x + 5y – z = 6 3x – y – 2z = 4 3x – 2y + 3z = 13 ( x + 5y + 6 x + 5y – 6 –x – 5y + 6 –x – 5y – 6 )

OpenStudy (anonymous):

from 1st... x+5y-z=6 x+5y-6=z so ....z=x+5y-6

OpenStudy (anonymous):

:) thanks Eliminate the “z” variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 (4x – 3y = –14 4x + y = –14 6x + 5y = 13 6x – 5y = –13 )